# GraphGDel: Constructing and Learning Graph Representations of Genome-Scale Metabolic Models for Growth-Coupled Gene Deletion Prediction

### v6

@@ V Discussion and Conclusion
In the computational experiments, we compared GraphGDel with two baseline methods (DNN and DeepGdel) on three constraint-based metabolic models of different scales. The results demonstrate the effectiveness of GraphGDel for predicting gene deletion strategies across models of varying sizes, consistently outperforming both baselines across all five evaluation metrics. Specifically, the overall accuracy of GraphGDel for e_coli_core, iMM904, and iML1515 was 77.35%, 85.54%, and 73.22%, respectively, representing improvements over the DNN baseline (14.04%, 16.26%, and 13.18% higher) and the DeepGdel baseline (6.17%, 4.96%, and 5.31% higher).
@@
A key innovation of GraphGDel is the explicit incorporation of graph structure information through the Graph-M module, addressing a limitation of previous approaches that relied solely on sequential data. By leveraging metabolic network topology, GraphGDel captures relationships between metabolites that are not apparent from their sequential representations. While this work uses LSTM autoencoders and Graph Convolutional Networks as the default architecture to demonstrate and benchmark the framework’s feasibility, its modular design allows for the integration of more advanced models. For example, one potential improvement is to incorporate recent large language models specialized for molecular or protein sequences into the learning modules of GraphGDel.
In future work, we will explore enhancing the graph representation construction pipeline to incorporate additional types of model elements, such as reactions and genes. Additionally, we will benchmark advanced deep learning architectures for metabolite and gene representation learning within the proposed framework across metabolic models from a broader range of species.
@@ VI ACKNOWLEDGMENTS
This work was supported in part by JSPS, KAKENHI under Grant 23K20386, and JST SPRING, Grant Number JPMJSP2110.

### v7

@@ V Discussion and Conclusion
In the computational experiments, we compared GraphGDel with three baseline methods (DNN, DeepGdel, and NMA) on three constraint-based metabolic models of different scales. The results demonstrate the effectiveness of GraphGDel for predicting gene deletion strategies across models of varying sizes, consistently outperforming all baseline methods across all five evaluation metrics. Specifically, the overall accuracy of GraphGDel for e_coli_core, iMM904, and iML1515 was 77.35%, 85.54%, and 73.22%, respectively, representing improvements over the DNN baseline (14.04%, 16.26%, and 13.18% higher), the DeepGdel baseline (6.17%, 4.96%, and 5.31% higher), and the NMA baseline (5.10%, 4.36%, and 4.70% higher).
@@
A key innovation of GraphGDel is the explicit incorporation of graph structure information through the Graph-M module, addressing a limitation of previous approaches that relied solely on sequential data. By leveraging metabolic network topology, GraphGDel captures relationships between metabolites that are not apparent from their sequential representations. While this work uses LSTM autoencoders and Graph Convolutional Networks as the default architecture to demonstrate and benchmark the framework’s feasibility, its modular design allows for the integration of more advanced models. For example, one potential improvement is to incorporate recent large language models specialized for molecular or protein sequences into the learning modules of GraphGDel. A limitation of this study is that the ground-truth labels are derived from constraint-based simulations curated in the MetNetComp database. Validation against independent biological experiments or alternative re-simulation frameworks was beyond the scope of the current work, and extending the evaluation to such settings remains an important direction for future research. Furthermore, the current framework does not explicitly provide pathway-level or graph-level explanations of its predictions. Incorporating interpretability techniques, such as attention mechanisms or graph attribution methods to highlight influential subgraphs or pathways, represents an important direction for future work. In this study, we focused on maximal gene deletion strategies. In the reference database, minimal deletion strategies are derived from maximal deletion sets; therefore, maximal deletions contain richer structural information and provide a more comprehensive representation of the feasible perturbation space. Importantly, the proposed framework is not inherently restricted to maximal deletions. Because it operates on general gene perturbation patterns, it can be readily extended to minimal or experimentally constrained intervention strategies. Exploring such practically constrained settings will be an important direction for future work.
In future work, we plan to enhance the graph representation construction pipeline by incorporating additional types of model elements, such as reactions and genes. We also aim to benchmark more advanced deep learning architectures for metabolite and gene representation learning across metabolic models from a broader range of species.
@@ VI ACKNOWLEDGMENTS
This work was supported in part by JSPS, KAKENHI under Grant 23K20386, 26K02972, and JST SPRING, Grant Number JPMJSP2110.